Mike is cutting the grass using a human-powered lawn mower. He pushes the mower with a force of 45 N directed at an angle of 41 N below the horizontal direction. Calculate the work that Mike does on the mower in pushing it 9.1 m across the hard.
1
Expert's answer
2020-04-13T09:58:20-0400



Initial data are shown on the figure. The projection of the applied to the lawn mower force onto the ground surface is:

"F_{pr} = 45 N \cdot \cos{41\degree} \approx 45N\cdot 0.75 = 33.75N.\\"

By definition, the work of constant force is the product of this force and the distance, on which this work was actihg on the body. Thus:

"A = F_{pr}\cdot9.1m = 33.75N\cdot9.1m = 307.125 J."


Answer. The work is A = 307.125J.